# Policy Mirror Descent with Lookahead

### v2

@@ 1 Introduction
Policy Mirror Descent (PMD) is a general class of algorithms for solving reinforcement learning (RL) problems. Motivated by the surge of interest in understanding popular Policy Gradient (PG) methods, PMD has been recently investigated in a line of works (Lan, 2023; Xiao, 2022; Li et al., 2023; Johnson et al., 2023). Notably, PMD encompasses several PG methods as particular cases via its flexible mirror map. A prominent example is the celebrated Natural PG method. PMD has also close connections to state-of-the-art methods such as TRPO and PPO (Tomar et al., 2022) which have achieved widespread empirical success (Schulman et al., 2015, 2017), including most recently in fine-tuning Large Language Models via RL from human feedback (Ouyang et al., 2022). Interestingly, PMD has also been inspired by one of the most fundamental algorithms in RL: Policy Iteration (PI). While PI alternates between policy evaluation and policy improvement, PMD regularizes the latter step to address the instability issue of PI with inexact policy evaluation.
Policy Iteration and its variants have been extensively studied in the literature, see e.g. Bertsekas and Tsitsiklis (1996); Munos (2003); Scherrer (2013). In particular, PI has been studied in conjunction with the lookahead mechanism (Efroni et al., 2018a), i.e. using multi-step greedy policy improvement instead of single-step greedy policies. Intuitively, the idea is that the application of the Bellman operator multiple times before computing a greedy policy leads to a more accurate approximation of the optimal value function. Implemented via Monte Carlo Tree Search (MCTS), multi-step greedy policy improvement [^1] Note that this is different from $n$-step return approaches which are rather used for policy evaluation (Sutton and Barto, 2018). contributed to the empirical success of some of the most impressive applications of RL including AlphaGo, AlphaZero and MuZero (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2019). Besides this practical success, a body of work (Efroni et al., 2018a, b; Tomar et al., 2020; Winnicki and Srikant, 2023; Winnicki et al., 2021) has investigated the role of lookahead in improving the performance of RL algorithms, reporting a convergence rate speed-up when enhancing PI with $h$-step greedy policy improvement with a reasonable computational overhead.
@@
  • We propose a novel class of algorithms called $h$-PMD enhancing PMD with multi-step greedy policy updates where $h$ is the depth of the lookahead. This class collapses to standard PMD when $h=1\,.$ When the stepsize parameter goes to infinity, we recover the PI algorithm with multiple-step greedy policy improvement previously analyzed in Efroni et al. (2018a). When solving a Markov Decision Process problem with $h$-PMD, we show that the value suboptimality gap converges to zero geometrically with a contraction factor of $\gamma^{h}$. This rate is faster than the standard convergence rate of PI and the similar rate for PMD which was recently established in Johnson et al. (2023) for $h\geq 2$. This rate improvement requires the computation of a $h$-step lookahead value function at each iteration which can be performed using planning methods such as tree-search. These results for exact $h$-PMD are exposed in Sec. 4, when value functions can be evaluated exactly.
@@
  • We perform simulations to verify our theoretical findings empirically on the standard DeepSea RL environment from Deep Mind’s bsuite (Osband et al., 2020). Our experiments illustrate the convergence rate improvement of $h$-PMD with increasing lookahead depth $h$ in both the exact and inexact settings.
@@ 2 Preliminaries
Value functions and optimal policy. For any policy $\pi\in\Pi$, we define the state value function $V^{\pi}:\mathcal{S}\to{{\mathbb{R}}}$ for every $s\in\mathcal{S}$ by $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]\,.$ The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\to{{\mathbb{R}}}$ can be similarly defined for every $s\in\mathcal{S},a\in\mathcal{A}$ by $Q^{\pi}(s,a)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$ where $\operatorname{\mathbb{E}}_{\pi}$ is the expectation over the state-action Markov chain $(s_{t},a_{t})$ induced by the MDP and the policy $\pi$ generating actions. The goal is to find a policy $\pi$ maximizing $V^{\pi}$. A classic result shows that there exists an optimal deterministic policy $\pi^{\star}\in\Pi$ maximizing $V^{\pi}$ simultaneously for all the states (Puterman, 2014). In this work, we focus on searching for an $\epsilon$-optimal policy, i.e. a policy $\pi$ satisfying $\|V^{\star}-V^{\pi}\|_{\infty}\leq\epsilon\,$ where $V^{\star}:=V^{\pi^{\star}}\,.$
@@ 3.1 PMD and its Connection to PI
Policy Gradient methods consist in performing gradient ascent updates with respect to an objective $\mathbb{E}_{\rho}[V^{\pi}(s)]$ w.r.t. the policy $\pi$ parametrized by its values $\pi(a|s)$ for $s,a\in\mathcal{S}\times\mathcal{A}$ in the case of a tabular policy parametrization. Interestingly, policy gradient methods have been recently shown to be closely connected to a class of Policy Mirror Descent algorithms by using dynamical reweighting of the Bregman divergence with discounted visitation distribution coefficients in a classical mirror descent update rule using policy gradients. We refer the reader to section 4 in Xiao (2022) or section 3.1 in Johnson et al. (2023) for a detailed exposition. The resulting PMD update rule is given by
@@
where we suppose throughout this paper that the function $\phi$ is of Legendre type, i.e. strictly convex and essentially smooth in the relative interior of $\operatorname{dom}\phi$ (see Rockafellar (1970), section 26). The mirror map choice gives rise to a large class of algorithms. Two special cases are noteworthy: (a) When $\phi$ is the squared $2$-norm, the Bregman divergence is the squared Euclidean distance and the resulting algorithm is a projected $Q$-ascent and (b) when $\phi$ is the negative entropy, the corresponding Bregman divergence is the Kullback-Leibler (KL) divergence and (PMD) becomes the popular Natural Policy Gradient algorithm (see e.g. Xiao (2022), section 4 for details).
@@
Connection to Policy Iteration. As previously noticed in the literature (Xiao, 2022; Johnson et al., 2023), PMD can be interpreted as a ‘soft’ PI. Indeed, taking infinite stepsizes $\eta_{k}$ in PMD (or a null Bregman divergence) immediately leads to the following update rule:
@@ 3.2 Policy Iteration with Lookahead
As discussed in Efroni et al. (2018a), Policy Iteration can be generalized by performing $h\geq 1$ greedy policy improvement steps at each iteration instead of a single step. The resulting $h$-PI update rule is:
@@
where $h\in\mathbb{N}\setminus\{0\}\,.$ The $h$-greedy policy $\pi^{k+1}$ w.r.t $V^{\pi_{k}}$ selects the first optimal action of a non-stationary $h$-horizon optimal control problem. It can also be seen as the $1$-step greedy policy w.r.t. ${\mathcal{T}}^{h-1}V^{\pi_{k}}\,.$ In the rest of this paper, we will denote by $\mathcal{G}_{h}(V)$ the set of $h$-step greedy policies w.r.t. any value function $V\in{{\mathbb{R}}}^{S}$, i.e. $\mathcal{G}_{h}(V):=\operatorname{argmax}_{\pi\in\Pi}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V=\{\pi\in\Pi:{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V={\mathcal{T}}^{h}V\}$. Notice that $\mathcal{G}_{1}(V)=\mathcal{G}(V)$. Overall, the $h$-PI method alternates between finding a $h$-greedy policy and estimating the value of this policy. Interestingly, similarly to PI, a $h$-greedy policy guarantees monotonic improvement, i.e. $V^{\pi^{\prime}}\geq V^{\pi}$ component-wise for any $\pi\in\Pi,\pi^{\prime}\in\mathcal{G}_{h}(V^{\pi})\,.$ Moreover, since ${\mathcal{T}}^{h}$ is a $\gamma^{h}$-contraction, it can be shown that the sequence $(\|V^{\star}-V^{\pi_{k}}\|_{\infty})_{k}$ is contracting with coefficient $\gamma^{h}\,.$ See section 3 in Efroni et al. (2018a) for further details about $h$-PI.
@@
While the iteration complexity always improves with larger depth $h$ (see Theorem 3 in Efroni et al. (2018a)), each iteration becomes more computationally demanding. As mentioned in Efroni et al. (2019), when the model is known, the $h$-greedy policy can be computed with Dynamic Programming (DP) in linear time in the depth $h$.
Another possibility is to implement a tree-search of depth $h$ starting from the root state $s$ to compute the $h$-greedy policy in deterministic MDPs. We refer to Efroni et al. (2019) for further details. Sampling-based tree search methods such as Monte Carlo Tree Search (MCTS) (see Browne et al. (2012)) can be used to circumvent the exponential complexity in the depth $h$ of tree search methods.
@@ 4.1 $h$-PMD: Using $h$-Step Greedy Updates
Connection to AlphaZero. Before switching gears to the analysis, we point out an interesting connection between $h$-PMD and the successful class of AlphaZero algorithms Silver et al. (2017, 2018); Schrittwieser et al. (2019) which are based on heuristics.
It has been shown in Grill et al. (2020b) that AlphaZero can be seen as a regularized policy optimization algorithm, drawing a connection to the standard PMD algorithm (with $h=1$). We argue that AlphaZero is even more naturally connected to our $h$-PMD algorithm. Indeed, the $Q$ values in (Grill et al., 2020b, Eq. (7) p. 3) correspond more closely to the lookahead action value function $Q_{h}^{\pi}$ approximated via tree search instead of the standard $Q^{\pi}$ function with $h=1$. This connection clearly delineates the dependence on the lookahead depth (or tree search depth) $h$ which was not clear in Grill et al. (2020b). We have implemented a version of our $h$-PMD algorithm with Deep Mind’s MCTS implementation (see section D.6 and the code provided for details). Conducting further experiments to compare our algorithm to AlphaZero on similar large scale settings would be interesting. We are not aware of any theoretical convergence guarantee for AlphaZero which relies on many heuristics. In contrast, our $h$-PMD algorithm enjoys theoretical guarantees that we establish in the next sections.
@@ 4.2 Convergence Analysis for Exact $h$-PMD
Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. (2023). In this section, we establish a linear convergence rate for the subobtimality value function gap of the exact $h$-PMD algorithm with a contraction factor of $\gamma^{h}$ where $h$ is the lookahead depth. We assume for now that the value function $V^{\pi_{k}}$ and a greedy policy $\pi_{k+1}$ in ($h$-PMD) can be computed exactly. These assumptions will be relaxed in the next section dealing with inexact $h$-PMD.
@@
  • Compared to Johnson et al. (2023), our new algorithm achieves a faster $\gamma^{h}$-rate where $h$ is the depth of the lookahead. It should be noted that the approach in Johnson et al. (2023) is claimed to be optimal over all methods that are guaranteed to increase the probability of the greedy action at each timestep. Adding lookahead to the policy improvement step circumvents this restriction.
  • Unlike prior work (Xiao, 2022; Lan, 2023) featuring distribution mismatch coefficients which can scale with the size of the state space, our convergence rate does not depend on any other instance-dependent quantities. This is thanks to the analysis which circumvents the use of the performance difference lemma similarly to Johnson et al. (2023).
  • Choosing $c_{t}=\gamma^{2h(t+1)}c_{0}$ for a positive constant $c_{0}$ for every integer $t$, the bound becomes (after bounding the geometric sum): $\lVert V^{\star}-V^{\pi_{k}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{c_{0}}{(1-\gamma)^{2}}\right)$. As $c_{0}\rightarrow 0$ we recover the linear convergence result of $h$-PI established in Efroni et al. (2018a).
@@
  • We defer the proof of Theorem 4.1 to Appendix A. Our proof highlights the relationship between $h$-PMD and $h$-PI through the explicit use of Bellman operators. Notice that even in the particular case of $h=1$, our proof is more compact compared to the one in Johnson et al. (2023) (see sec. 6 and Lemma A.1 to A.3 therein for comparison) using our convenient notations.
@@ 5.1 Monte Carlo $h$-Greedy Policy Evaluation
Notice that actions are exhaustively selected in our procedure like in a planning method. Using bandit ideas to guide Monte Carlo planning (see e.g. Kocsis and Szepesvári (2006a)) would be interesting to investigate for more efficiency. We leave the investigation of such selective action sampling and exploration for future work. In practice, lookahead policies are often computed using tree search techniques (Efroni et al., 2018a).

### v3

@@ 1 Introduction
Policy Mirror Descent (PMD) is a general class of algorithms for solving reinforcement learning (RL) problems. Motivated by the surge of interest in understanding popular Policy Gradient (PG) methods, PMD has been recently investigated in a line of works [25, 53, 29, 20]. Notably, PMD encompasses several PG methods as particular cases via its flexible mirror map. A prominent example is the celebrated Natural PG method. PMD has also close connections to state-of-the-art methods such as TRPO and PPO [50] which have achieved widespread empirical success [41, 42], including most recently in fine-tuning Large Language Models via RL from human feedback [35]. Interestingly, PMD has also been inspired by one of the most fundamental algorithms in RL: Policy Iteration (PI). While PI alternates between policy evaluation and policy improvement, PMD regularizes the latter step to address the instability issue of PI with inexact policy evaluation.
Policy Iteration and its variants have been extensively studied in the literature, see e.g. [5, 32, 39]. In particular, PI has been studied in conjunction with the lookahead mechanism [12], i.e. using multi-step greedy policy improvement instead of single-step greedy policies. Intuitively, the idea is that the application of the Bellman operator multiple times before computing a greedy policy leads to a more accurate approximation of the optimal value function. Implemented via Monte Carlo Tree Search (MCTS), multi-step greedy policy improvement [^1] Note that this is different from $n$-step return approaches which are rather used for policy evaluation [47]. contributed to the empirical success of some of the most impressive applications of RL including AlphaGo, AlphaZero and MuZero [44, 45, 46, 40]. Besides this practical success, a body of work [12, 13, 49, 51, 52] has investigated the role of lookahead in improving the performance of RL algorithms, reporting a convergence rate speed-up when enhancing PI with $h$-step greedy policy improvement with a reasonable computational overhead.
@@
  • We propose a novel class of algorithms called $h$-PMD enhancing PMD with multi-step greedy policy updates where $h$ is the depth of the lookahead. This class collapses to standard PMD when $h=1\,.$ When the stepsize parameter goes to infinity, we recover the PI algorithm with multiple-step greedy policy improvement previously analyzed in [12]. When solving a Markov Decision Process problem with $h$-PMD, we show that the value suboptimality gap converges to zero geometrically with a contraction factor of $\gamma^{h}$. This rate is faster than the standard convergence rate of PI and the similar rate for PMD which was recently established in [20] for $h\geq 2$. This rate improvement requires the computation of a $h$-step lookahead value function at each iteration which can be performed using planning methods such as tree-search. These results for exact $h$-PMD are exposed in Sec. 4, when value functions can be evaluated exactly.
@@
  • We perform simulations to verify our theoretical findings empirically on the standard DeepSea RL environment from Deep Mind’s bsuite [34]. Our experiments illustrate the convergence rate improvement of $h$-PMD with increasing lookahead depth $h$ in both the exact and inexact settings.
@@ 2 Preliminaries
Value functions and optimal policy. For any policy $\pi\in\Pi$, we define the state value function $V^{\pi}:\mathcal{S}\to{{\mathbb{R}}}$ for every $s\in\mathcal{S}$ by $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]\,.$ The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\to{{\mathbb{R}}}$ can be similarly defined for every $s\in\mathcal{S},a\in\mathcal{A}$ by $Q^{\pi}(s,a)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$ where $\operatorname{\mathbb{E}}_{\pi}$ is the expectation over the state-action Markov chain $(s_{t},a_{t})$ induced by the MDP and the policy $\pi$ generating actions. The goal is to find a policy $\pi$ maximizing $V^{\pi}$. A classic result shows that there exists an optimal deterministic policy $\pi^{\star}\in\Pi$ maximizing $V^{\pi}$ simultaneously for all the states [36]. In this work, we focus on searching for an $\epsilon$-optimal policy, i.e. a policy $\pi$ satisfying $\|V^{\star}-V^{\pi}\|_{\infty}\leq\epsilon\,$ where $V^{\star}:=V^{\pi^{\star}}\,.$
@@ 3.1 PMD and its Connection to PI
Policy Gradient methods consist in performing gradient ascent updates with respect to an objective $\mathbb{E}_{\rho}[V^{\pi}(s)]$ w.r.t. the policy $\pi$ parametrized by its values $\pi(a|s)$ for $s,a\in\mathcal{S}\times\mathcal{A}$ in the case of a tabular policy parametrization. Interestingly, policy gradient methods have been recently shown to be closely connected to a class of Policy Mirror Descent algorithms by using dynamical reweighting of the Bregman divergence with discounted visitation distribution coefficients in a classical mirror descent update rule using policy gradients. We refer the reader to section 4 in [53] or section 3.1 in [20] for a detailed exposition. The resulting PMD update rule is given by
@@
where we suppose throughout this paper that the function $\phi$ is of Legendre type, i.e. strictly convex and essentially smooth in the relative interior of $\operatorname{dom}\phi$ (see [37], section 26). The mirror map choice gives rise to a large class of algorithms. Two special cases are noteworthy: (a) When $\phi$ is the squared $2$-norm, the Bregman divergence is the squared Euclidean distance and the resulting algorithm is a projected $Q$-ascent and (b) when $\phi$ is the negative entropy, the corresponding Bregman divergence is the Kullback-Leibler (KL) divergence and (PMD) becomes the popular Natural Policy Gradient algorithm (see e.g. [53], section 4 for details).
@@
Connection to Policy Iteration. As previously noticed in the literature [53, 20], PMD can be interpreted as a ‘soft’ PI. Indeed, taking infinite stepsizes $\eta_{k}$ in PMD (or a null Bregman divergence) immediately leads to the following update rule:
@@ 3.2 Policy Iteration with Lookahead
As discussed in [12], Policy Iteration can be generalized by performing $h\geq 1$ greedy policy improvement steps at each iteration instead of a single step. The resulting $h$-PI update rule is:
@@
where $h\in\mathbb{N}\setminus\{0\}\,.$ The $h$-greedy policy $\pi^{k+1}$ w.r.t $V^{\pi_{k}}$ selects the first optimal action of a non-stationary $h$-horizon optimal control problem. It can also be seen as the $1$-step greedy policy w.r.t. ${\mathcal{T}}^{h-1}V^{\pi_{k}}\,.$ In the rest of this paper, we will denote by $\mathcal{G}_{h}(V)$ the set of $h$-step greedy policies w.r.t. any value function $V\in{{\mathbb{R}}}^{S}$, i.e. $\mathcal{G}_{h}(V):=\operatorname{argmax}_{\pi\in\Pi}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V=\{\pi\in\Pi:{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V={\mathcal{T}}^{h}V\}$. Notice that $\mathcal{G}_{1}(V)=\mathcal{G}(V)$. Overall, the $h$-PI method alternates between finding a $h$-greedy policy and estimating the value of this policy. Interestingly, similarly to PI, a $h$-greedy policy guarantees monotonic improvement, i.e. $V^{\pi^{\prime}}\geq V^{\pi}$ component-wise for any $\pi\in\Pi,\pi^{\prime}\in\mathcal{G}_{h}(V^{\pi})\,.$ Moreover, since ${\mathcal{T}}^{h}$ is a $\gamma^{h}$-contraction, it can be shown that the sequence $(\|V^{\star}-V^{\pi_{k}}\|_{\infty})_{k}$ is contracting with coefficient $\gamma^{h}\,.$ See section 3 in [12] for further details about $h$-PI.
@@
While the iteration complexity always improves with larger depth $h$ (see Theorem 3 in [12]), each iteration becomes more computationally demanding. As mentioned in [14], when the model is known, the $h$-greedy policy can be computed with Dynamic Programming (DP) in linear time in the depth $h$.
Another possibility is to implement a tree-search of depth $h$ starting from the root state $s$ to compute the $h$-greedy policy in deterministic MDPs. We refer to [14] for further details. Sampling-based tree search methods such as Monte Carlo Tree Search (MCTS) (see [9]) can be used to circumvent the exponential complexity in the depth $h$ of tree search methods.
@@ 4.1 $h$-PMD: Using $h$-Step Greedy Updates
Connection to AlphaZero. Before switching gears to the analysis, we point out an interesting connection between $h$-PMD and the successful class of AlphaZero algorithms [45, 46, 40] which are based on heuristics.
It has been shown in [17] that AlphaZero can be seen as a regularized policy optimization algorithm, drawing a connection to the standard PMD algorithm (with $h=1$). We argue that AlphaZero is even more naturally connected to our $h$-PMD algorithm. Indeed, the $Q$ values in [17, Eq. (7) p. 3] correspond more closely to the lookahead action value function $Q_{h}^{\pi}$ approximated via tree search instead of the standard $Q^{\pi}$ function with $h=1$. This connection clearly delineates the dependence on the lookahead depth (or tree search depth) $h$ which was not clear in [17]. We have implemented a version of our $h$-PMD algorithm with Deep Mind’s MCTS implementation (see section D.6 and the code provided for details). Conducting further experiments to compare our algorithm to AlphaZero on similar large scale settings would be interesting. We are not aware of any theoretical convergence guarantee for AlphaZero which relies on many heuristics. In contrast, our $h$-PMD algorithm enjoys theoretical guarantees that we establish in the next sections.
@@ 4.2 Convergence Analysis for Exact $h$-PMD
Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in [20]. In this section, we establish a linear convergence rate for the subobtimality value function gap of the exact $h$-PMD algorithm with a contraction factor of $\gamma^{h}$ where $h$ is the lookahead depth. We assume for now that the value function $V^{\pi_{k}}$ and a greedy policy $\pi_{k+1}$ in ($h$-PMD) can be computed exactly. These assumptions will be relaxed in the next section dealing with inexact $h$-PMD.
@@
  • Compared to [20], our new algorithm achieves a faster $\gamma^{h}$-rate where $h$ is the depth of the lookahead. It should be noted that the approach in [20] is claimed to be optimal over all methods that are guaranteed to increase the probability of the greedy action at each timestep. Adding lookahead to the policy improvement step circumvents this restriction.
  • Unlike prior work [53, 25] featuring distribution mismatch coefficients which can scale with the size of the state space, our convergence rate does not depend on any other instance-dependent quantities. This is thanks to the analysis which circumvents the use of the performance difference lemma similarly to [20].
  • Choosing $c_{t}=\gamma^{2h(t+1)}c_{0}$ for a positive constant $c_{0}$ for every integer $t$, the bound becomes (after bounding the geometric sum): $\lVert V^{\star}-V^{\pi_{k}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{c_{0}}{(1-\gamma)^{2}}\right)$. As $c_{0}\rightarrow 0$ we recover the linear convergence result of $h$-PI established in [12].
@@
  • We defer the proof of Theorem 4.1 to Appendix A. Our proof highlights the relationship between $h$-PMD and $h$-PI through the explicit use of Bellman operators. Notice that even in the particular case of $h=1$, our proof is more compact compared to the one in [20] (see sec. 6 and Lemma A.1 to A.3 therein for comparison) using our convenient notations.
@@ 5.1 Monte Carlo $h$-Greedy Policy Evaluation
Notice that actions are exhaustively selected in our procedure like in a planning method. Using bandit ideas to guide Monte Carlo planning (see e.g. [23]) would be interesting to investigate for more efficiency. We leave the investigation of such selective action sampling and exploration for future work. In practice, lookahead policies are often computed using tree search techniques [12].